OpenStudy (anonymous):

Multiply 3/8x and 2/5x ... both 3/8 and 2/5 are exponents of x. plz help I think the answer is 5/13x but not sure.

OpenStudy (anonymous):

Do you mean: \[x^{\frac{3}{8}} * x^{\frac{2}{5}}\]

OpenStudy (anonymous):

yes.

OpenStudy (lgbasallote):

5/13 is not quite right

OpenStudy (anonymous):

okay, idk what I did wrong?

OpenStudy (lgbasallote):

@Wired should you or should i

OpenStudy (anonymous):

You can go for it :)

OpenStudy (lgbasallote):

\[\huge x^{3/8} \cdot x^{2/5} \implies x^{3/8 + 2/5}\] agree?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

so 31/40x thank u again :)

OpenStudy (lgbasallote):

yup there you go